how the displacement of a vertically thrown ball varies with time
1
Expert's answer
2019-11-06T09:11:45-0500

               If the body is thrown vertically up in the presence of an initial speed vy1 = 0 (Fig.1).    It will move equally slowly with an acceleration  equal  to  a= −g = − 9,81  m/s2

 




                                                                     Fig.1

     When the body reaches a height equal to  Hmax   then it has speed  vy1 = 0  and acceleration g. It follows that the body will not be able to stay at this height, so it will go into a state of free fall.

     That is, a body thrown upwards is an equally slow motion, in which, after reaching  Hmax  signs of movement are  reversed.

     The basic formulas for calculating the motion of a body thrown vertically up are

Speed at any given time  t                  

                            v  = v0  - gt ,   m/s.

Moving at any time                                                 

H= v0t - gt2/2, m.

Coordinate at any time t                                             

y=y0 +H= v0t - gt2/2, m